Sign in
Please select an account to continue using cracku.in
↓ →
The probability that a certain electronic component fails when first used is 0.10. If it does not fail immediately, then the probability that it lasts for one year is 0.99. The probability that a new component will last for one year is
The probability that the component fails when used for the first time = 0.1 (or)
The probability that the component does not fail when used for the first time = 1 - 0.1 = 0.9
The probability that the component lasts for one year = 0.99
Hence, the probability that the component will last for one year = 0.9 x 0.99 = 0.891
Hence, option A is the correct answer.
Was this solution helpful?
Create a FREE account and get:
Book Free CAT Mentorship
Get personalized CAT strategy from a 99%iler
500+ students mentored
OTP Verification
Enter the 6-digit code sent to your phone
Booking Summary
Enter OTP
Didn't receive the OTP?
Start your IIM journey with the right preparation and crack CAT 2026.
Educational materials for CAT preparation